4. The catch-up effect Consider the economies of Sporon and Gribinez, both of which produce gaggles of gop using only tools and workers. Suppose that, during the course of 40 years, the level of physical capital per worker rises by 5 tools per worker in each economy, but the size of each labor force remains the same Complete the following tables by entering productivity (in terms of output per worker) for each economy in 2016 and 2056 Sporon Physical Capita (Tools per worker) 7 12 Output (Gaggles of gop) 1,800 2,160 bor Force Productivity (Gaggles per worker) Year 2016 2056 (Workers) 30 30 60 72 Gribinez Physical Capita abor Force Output (Gaggles of gop) 900 1,620 Productivity (Gaggles per worker) Year 2016 2056 (Tools per worker) (Workers) 30 30 4 30 54 Initially, the number of tools per worker was higher in Sporon than in Gribinez. From 2016 to 2056, capital per worker rises by 5 units in each country. The 5-unit change in capital per worker causes productivity in Sporon to rise by a larger ▼ amount than productivity in Gribinez. This illustrates the effect.Gribinez Physical Capital Labor Forceutnut Productivity Year (Tools per worker) (Worker human capital op(Gaggles per worker) 2016 2056 4 30 30 catch-up natural resources technology 30 54 Initially, the number of tools per worker was units in each country. The 5-unit change in productivity in Gribinez. This illustrates the an in Gribinez. From 2016 to 2056, capital per worker rises by 5 auses productivity in Sporon to rise by a _largeramount than effect.
Answer #1

For sporon:

Productivity : 2016, output/ Labor force = 1800/30= 60

In year 2056, productivity = 2160/30 = 72

For gribinez:

2016, productivity = 900/30 = 30

2056, productivity = 1620/30 = 54

Now productivity change in % for Sporon =( 72-60/60)*100

= 20 %

For Gribinez: (54-30/30)*100 = 80%

Thus larger rise in productivity for Gribinez

Now this effect is known as Catch up effect, which is defined as convergence in development economics, that over the time poorer economies tend to grow faster than the economies with a higher per capita income and gradually reach similar high levels of per capita income.